### v1

# Two Causally Related Needles in a Video Haystack

Abstract
Evaluating the video understanding capabilities of Video-Language Models (VLMs) remains a significant challenge. We propose a long-context video understanding benchmark, Causal2Needles, that assesses two crucial abilities insufficiently evaluated by existing benchmarks: (1) the ability to extract information from two separate locations in a long video and understand them jointly, and (2) the ability to model the world in terms of cause and effect in human behaviors. Specifically, Causal2Needles introduces 2-needle questions, which require extracting information from both the cause and effect human-behavior events in a long video and the associated narration text. To prevent textual bias, these questions comprise two complementary formats: one asking to identify the video clip containing the answer, and one asking for the textual description of an unrelated visual detail from that video clip. Our experiments reveal that models excelling in pre-existing benchmarks struggle with 2-needle visual grounding, and the model performance is negatively correlated with the distance between the two needles. These findings highlight critical limitations in current VLMs. The dataset is available at: https://huggingface.co/datasets/causal2needles/Causal2Needles
1 Introduction
On many popular benchmarks (Hendrycks et al., 2020; Cobbe et al., 2021; Srivastava et al., 2022; Chen et al., 2021), recent AI systems achieved performance comparable to humans. However, real-world observations suggest that there is still a significant gap between AI and human capabilities (West et al., 2023; Frieder et al., 2023; Amirizaniani et al., 2024). This apparent contradiction suggests that (1) there is much overfitting to popular benchmarks such as GSM8K (Mirzadeh et al., 2024) through mechanisms like data contamination (Singh et al., 2024), and (2) existing benchmarks may not fully reflect differences between machine intelligence and human intelligence (Dziri et al., 2023; Wu et al., 2024). As a result, the development of sophisticated benchmarks that critically evaluate model capabilities has become a high priority for AI research.
Within the context of long video understanding, we investigate two crucial limitations of existing benchmarks for Video-Language Models (VLMs). First, benchmarks evaluating information extraction from a single location do not fully reflect the long video understanding ability of VLMs. A popular type of evaluation for long-context models is called the “needle in a haystack” problem (Kamradt, 2023; Wang et al., 2024b), where the “needle” represents the information to be extracted from the context. Some long video benchmarks feature 1-needle questions, requiring information from a single location in the video. However, research in NLP indicates that models excelling at 1-needle questions often experience significant performance drops when required to extract multiple needles or understand them jointly, revealing the limitations of 1-needle questions for assessing long-context understanding (Li et al., 2023; Vodrahalli et al., 2024; Yang, 2024; Levy et al., 2024). However, multiple-needle questions remain rare in the multimodal setting.
Second, existing benchmarks offer an incomplete evaluation of whether VLMs possess an internal world model. A world model captures the underlying mechanisms governing object dynamics and human behaviors, enabling predictions of future events given an intervention (Forrester, 1971; Ha and Schmidhuber, 2018). However, existing evaluation of world models focus solely on object motion prediction while neglecting human behaviors (Guan et al., 2024a; Motamed et al., 2025; Kang et al., 2024).
To address these limitations, we propose a long-context video understanding benchmark, Causal2Needles, including 902 1-needle questions and 1,494 2-needle questions. It evaluates two key abilities of VLMs: (1) the ability to extract relevant information from two locations in long videos and jointly reason about them, and (2) the ability to model the world in terms of cause and effect of human behaviors.
To evaluate the first ability, we design 2-needle questions using a bridge entity, or the information shared between two causally related events. The identity of the bridge entity is purposely obscured in the 2-needle question, so that the second needle cannot be retrieved before the identity is resolved, which in turn requires the first needle to be retrieved. This mechanism forces the model to jointly understand the two needles in order to answer the question. As a concrete example, Fig. 1 shows a cause event and an effect event that share a bridge entity, “Superman’s death,” ambiguously referred to as “tragedy.” To answer part 1 of the question, the model must first retrieve the effect event, which reveals the tragedy is Superman’s death. Only after that, the model can answer part 2 by retrieving the clip showing the cause event. Suppose the bridge entity were explicitly stated, the question would become “how did Superman die, leading to his memorial service in Metropolis” and be answerable by retrieving the cause event directly. In that case, joint understanding of two needles would be unnecessary and the 2-needle question would degenerate to a 1-needle question.
To assess the second ability, we design our benchmark to test VLMs on identifying the causes of events across diverse, often fictional scenarios. Causal reasoning is fundamental for understanding how actions lead to outcomes, enabling models to go beyond surface patterns to grasp underlying mechanisms (Pearl et al., 2000). Causal2Needles evaluates the understanding of human behaviors from a causal perspective, complementing existing assessments of world models.
Figure 1: The logical solution process for the 2-needle questions of Causal2Needles. Each step involves an operation, with the input shown above the step and the output below the step. The question purposely refer to the bridge entity, “Superman’s death,” ambiguously as “tragedy.” As a result, one must first resolve the bridge entity using Part 1 before answering Part 2. This question design mandates joint understanding of both the cause and effect events. Note that the steps are necessary only in an information-processing sense. A VLM may adopt different steps.
Figure 2: The evaluation framework of Causal2Needles. To help models understand the storyline, we also feed the full textual narration into the model. Three types of questions are designed for each pair of causally related events.
A potential drawback of Causal2Needles is that the video clips could be difficult to understand without the narration, but the narration text could introduce textual bias that allows questions to be answered from text alone without video (Ko et al., 2023; Cores et al., 2024; Xiao et al., 2024). As mitigation, we introduce two complementary question formats. The first format, which we call the visual grounding question, requires the model to select the video clip corresponding to the event it needs to retrieve. The second format, which we call the image description question, requires the model to answer multiple-choice questions about the appearance of the retrieved video clip. Fig. 2 shows some examples. Visual grounding questions necessitate the understanding of video clips, but they may underestimate model performance as the format may be out-of-distribution (OOD) to most VLMs. In contrast, image description questions circumvent the OOD issue, but the VLM may benefit from general knowledge of the movie from pretraining data, such as the color of Lois Lane’s outfit, leading to overestimation of model abilities. By evenly splitting 2-needle questions between the two question formats (747 each), Causal2Needles establish a comprehensive benchmark.
Experiments reveal several important findings. First, models that perform well on 1-needle questions exhibit remarkable performance drops on 2-needle questions. Second, the distance between the two needles is negatively correlated with model performance. Taken together, these results suggest that joint understanding of two separate video locations is an important weakness of existing VLMs. We summarize the contributions of this paper as follows:
$\bullet$ We propose the Causal2Needles benchmark that contains 2-needle questions, generated from story videos. We devise a question creation strategy utilizing bridge entities to force the video-language models to jointly understand two video clips arbitrarily located in the context.
$\bullet$ We reveal significant weaknesses in current VLMs in the joint understanding and retrieval of two separate video clips, even though their 1-needle performances are high. The more distant the two clips are, the worse the models perform.
2 Related Work
2.1 Long Video Understanding Benchmarks
The ability to retrieve information from a specific location in a long video (often called the “needle in a haystack” task) is widely regarded as a key performance metric for the long-video understanding ability (Kamradt, 2023; Wang et al., 2024b). In Tab. 1, we compare recent long-video benchmarks based on needle-in-haystack tasks side by side to highlight the unique contribution of our Causal2Needles.
We focus exclusively on benchmarks involving multi-needle questions. EgoSchema (Mangalam et al., 2023) fail to assess specific model capabilities due to the lack of type-wise evaluation (e.g., 1-needle vs. 2-needle, reasoning vs. recognition), resulting in limited diagnostic precision. MLVU (Zhou et al., 2024) require models to retrieve visually distinct images that are artificially inserted into videos, which may allow for shortcut features due to the domain gap between the artificial needle and the original content. In contrast to datasets dominated by single-needle questions, MVBench (Li et al., 2024c) and TVBench (Cores et al., 2024) include a large proportion of temporal tracking questions that evaluate a model’s ability to track the state of an object across multiple temporal locations. A similar type of questions is also featured in LongVideoBench (Wu et al., 2025) (e.g., L2-SSS questions) and VideoMME (Fu et al., 2024) (e.g., Counting problems). These temporal tasks typically require only independent understanding of each needle. Moreover, the relationships between their needles often rely on surface-level visual cues, such as matching object colors or identities.
Table 1: A comparison of Causal2Needles with other multi-needle long video benchmarks. Causal2Needles is the only benchmark dedicated to needle-in-haystack problems (Diagnostic Precision) that requires joint understanding of the two needles and the identification of cause events from effect events.
Benchmark Video Length # QA Diagnostic Precision Needle Type Joint Understanding Causal Reasoning
EgoScheme (Mangalam et al., 2023) 180 s 5,000 Natural
MVBench (Li et al., 2024c) 16 s-40 s 4,000 Natural
TVBench (Cores et al., 2024) 16 s-40 s 2,525 Natural
MLVU (Zhou et al., 2024) 180 s-3600 s 3,102 Artificial
VideoMME (Fu et al., 2024) 1018s 2,700 Natural
LVB Wang et al. (2024a) 473 s 6,679 Natural
Causal2Needles (Ours) 438 s 2,396 Natural
Differing from the datasets above, our benchmark requires models to jointly understand both the cause and the effect—the understanding of one relies on the correct interpretation of the other. Furthermore, the relation between the two needles is causality rather than visual similarity. These designs establish a challenging and unique benchmark for long-video understanding.
2.2 World Model
There is an ongoing debate (Gurnee and Tegmark, 2023; Liu et al., 2025; Guan et al., 2024b; Hu et al., 2024; Kang et al., 2024; Motamed et al., 2025) over whether deep learning networks develop internal world models—whether they learn the underlying principles governing entities, which can be used to to predict future outcomes of possible interventions (Forrester, 1971; Ha and Schmidhuber, 2018).
However, this debate largely overlooks causality between events involving humans, which could be induced by factors like biology (e.g., heavy sweating causes thirst), psychology (e.g., hearing compliments makes people feel good), and social norms (e.g., providing good service leads to tipping). Although such causal questions occasionally appear in video understanding datasets, for example in the form of episodic reasoning in MVBench, it is usually difficult to isolate them and quantitatively measure this factor independently. In contrast, Causal2Needles is dedicated to evaluating reasoning over cause and effect in the context of human behaviors, which allows precise diagnosis of this capability and fills a gap in benchmarks of world models.
3 Causal2Needles: Dataset Construction
Causal2Needles is built on two video-language datasets, YMS (Dogan et al., 2018) and SyMoN (Sun et al., 2022). With a total of 192 fully annotated movie recap videos, the datasets offer a rich collection of human-behavior events from diverse movie genres. Each event consists of a narration sentence and its corresponding video clip. Details such as the distribution of movie themes and the temporal distance between causal events can be found in Appendix Sec. A.
Our primary focus is to generate challenging 2-needle questions. The generation process is based on a Large Language Model (LLM), ChatGPT-o1, and is decomposed into three steps. First, we extract event-level causal relationships (sentence pairs) from the narration of a video. Then, we rephrase each pair of sentences to explicitly establish a bridge entity between them. Finally, based on the bridge entity, we generate 2-needle questions requiring a joint understanding of both cause and effect events. Each step is guided by carefully designed prompts and illustrative examples (see Appendix Sec. F.1 for details), enabling the LLM to perform the generation incrementally and reliably. Representative examples of the generated questions are shown in Fig. 2. The three steps are elaborated in detail in Sec. 3.1 and 3.2.
3.1 Causal Relationships Extraction
We employ an LLM to extract causal relationships from narration by combining global and local event graphs, based on an event graph extraction method (Sun et al., 2024). Each event (a narration sentence) is represented as a node, and causal relationships as directed edges. The global graph is extracted from the complete narration of a video. The graph captures long-range causal relationships but may be incomplete, because LLM often overlook sentences in the middle section of a long context (Liu et al., 2024). To address this, we introduce a sliding-window approach that extracts local graphs from 15-sentence segments with a 5-sentence stride. These local graphs capture more detailed causal relationships with a shorter range. We merge the global and local graphs to obtain comprehensive and long-range causal relationships. To avoid superficial causal relationships resulting from temporal adjacency, we retain only those where the cause and effect are separated by at least three events. In Causal2Needles, the distance between cause and effect events ranges from 3 to 21. Appendix Sec. A has more details.
3.2 Generation of Two-Needle Questions
Generating 2-needle questions involves two steps after obtaining causal relationships: (1) rephrasing the cause and effect sentences to establish the bridge entity between them, and (2) generating a question that requires joint understanding of both cause and effect events.
Rephrasing the Cause and Effect Sentences.
The generation of 2-needle questions is based on a bridge entity that connects cause and effect events. However, even if two events have a causal relationship, their corresponding narration sentences may not contain an obvious bridge entity because the original narration does not explicitly emphasize causality. In light of this, we prompt a VLM (Gemini-1.5-pro) to rephrase the cause and effect sentences to explicitly establish the bridge entity. For example, the cause sentence, “As it dies, the creature stabs and kills the weakened Superman with one of its bone protrusions,” and the effect sentence, “A memorial is held for Superman in Metropolis,” are rephrased into sentences in Fig. 2. To maintain semantic consistency between the rephrased sentences and their corresponding scenes, we also input the scenes into the VLM as a constraint.
Generation of Visual Grounding Questions.
After obtaining causal sentence pairs with a clear bridge entity, we utilize an LLM to generate two parts of a visual grounding 2-needle question. We use the causal relationship in Fig. 2 to demonstrate this process.
First, we prompt the LLM to generate a Part 1 question that requires retrieving the effect event to identify the bridge entity. Specifically, we ask it to extract the bridge entity that establishes the causal relationship, such as “Superman’s death.” Next, we instruct the LLM to rephrase this bridge entity as an ambiguous term, such as “tragedy.” Finally, using this ambiguous term as a reference, we prompt the LLM to create a question that uses the effect sentence as context and the bridge entity as the answer. In this case, the final output is: “What tragedy was commemorated by the memorial service in Metropolis?”
Second, we prompt the LLM to generate a Part 2 question grounded on Part 1, which requires retrieving the cause event. Specifically, based on the ambiguous term “tragedy”, we instruct the LLM to generate a question whose answer appears exclusively in the cause sentence. The output is: “How did such a tragedy occur?”
Finally, combining the two parts with task instruction, we build a visual grounding 2-needle question: “What tragedy was commemorated by the memorial service in Metropolis, and how did such a tragedy occur? Select the scenes that contain the answers to the question.” To answer such questions, the model must jointly understand cause and effect events and ground the answers on video clips.
Generation of Image Description Questions.
Since the likely out-of-distribution visual grounding questions may underestimate the performance of VLMs, we also generate image description questions with text-option multiple-choice format. These questions share the same Part 1 as visual grounding questions, but modify Part 2 to ask about visual details of the cause event.
[…]
However, the generated questions may reveal the bridge entity or include excessive object details, allowing the model to locate the cause scene without truly understanding the Part 1 question. For instance, the initially generated question reveals the bridge entity “Superman’s death” and explicitly states the name “Doomsday”, reduce the difficulty of locating the correct scene.
To mitigate this issue, we prompt an LLM to obscure the bridge entity and the named entity being inquired about. We then obtain the final image description 2-needle question, which begins with the Part 1 question: “What tragedy was commemorated by the memorial service in Metropolis? When the tragedy that causes the memorial service occurs, what color were the energy blasts emanating from the creature? A. Yellow B. Blue C. Red D. White”
3.3 Generation of One-Needle Questions
With the extracted causal relationships, we use an LLM to generate simple 1-needle questions that ask for the cause of a given event. Specifically, we illustrate the process using the causal relationship in Fig. 2. First, we prompt the LLM to generate a question using the effect sentence as context, with the answer found exclusively in the cause sentence. The generated question is: “How did Superman die, leading to his memorial service in Metropolis?” Then, to mitigate textual bias from the input narration, we combine this question with the visual grounding instruction: “Select the scene that contains the answer to the question.”
3.4 Quality Evaluation of Generated Questions
We utilize two state-of-the-art models not involved in the question generation process, ChatGPT-4.1 and Gemini-2.0-flash, to evaluate the quality of questions in Causal2Needles. The evaluation experiments are conducted from three perspectives: (1) Bridge entity accuracy, or if the bridge entity used for 2-needle question generation correctly connects the corresponding cause and effect events. In addition to the ground-truth bridge entity, we also construct a random baseline by randomly matching the bridge entities and the cause-effect event pairs. This baseline helps to verify that the evaluation LLMs do not always output high accuracy. (2) Semantic coherence, or whether the question is consistent with the story and does not introduce hallucination or content that contradicts the story. We use a 5-point scale, where 1 is the lowest score and 5 the highest. In the random baseline, we shuffle the correspondence between the questions and the cause-effect event pairs. (3) Readability of the questions, as reflected by the naturalness, grammar, and clarity, on a 5-point scale. Due to the difficulty in writing unreadable questions, we do not construct the random baseline for readability. The prompts for question evaluation are in Appendix Sec. F.
Tab. 2 shows the evaluation results. The generated questions receive near-perfect scores on bridge entity accuracy and semantic coherence, and score highly on readability. Reassuringly, both random baselines score near the lowest possible, which is 0% for bridge entity accuracy and 1 for semantic coherence. These results indicate that the automatically generated questions have high quality and the evaluation processes is valid.
Table 2: Evaluation results of generated questions. ‘1-N’, ‘VG 2-N’ and ‘ID 2-N’ denote 1-needle questions, visual grounding 2-needle questions and image description 2-needle questions, respectively. Numbers in parentheses indicate the performance of random baselines.
Models Bridge Entity (Random Baseline) Semantic Coherence Readability
1-N VG 2-N ID 2-N 1-N VG 2-N ID 2-N
ChatGPT-4.1 95.6% (0.3%) 4.62 (1.12) 4.99 (1.10) 4.74 (1.13) 4.85 4.83 4.67
Gemini-2.0-flash 95.0% (3.8%) 4.66 (1.02) 4.96 (1.01) 4.83 (1.00) 4.18 4.69 4.25
Table 3: Quantitative results (accuracy, %) of VLMs on our benchmark. ‘Forward’ refers to inputting video scenes in chronological order, while ‘Reverse’ uses reverse order. ‘Avg’ denotes results averaged over both orders. Best scores are in bold. ‘ID’ denotes image description and ‘VG’ denotes visual grounding.
Models 1-Needle Questions VG 2-Needle Questions ID 2-Needle Questions
Forward Reverse Avg
Part 1 Part 2 Both Part 1 Part 2 Both Part 1 Part 2 Both
Proprietary Models
ChatGPT-4o 39.2 16.7 39.2 9.4 45.4 21.2 13.4 31.1 30.2 11.4 59.2
Gemini-1.5-pro 35.6 21.0 40.0 10.2 35.7 21.4 8.4 28.4 30.7 9.3 60.9
ChatGPT-4o-mini 33.4 17.4 22.9 5.0 32.4 11.9 5.2 24.9 17.4 5.1 52.3
Claude-3.5-sonnet 26.5 16.6 22.4 4.8 19.3 13.9 2.9 17.9 18.1 3.9 60.5
Open-source Models
LLaVA-OneVision-7B 18.0 4.6 14.7 0.0 17.0 5.6 0.1 10.8 10.2 0.1 28.3
LLaVA-Next-Video-7B 17.2 0.0 4.4 0.0 15.9 0.0 0.0 8.0 2.2 0.0 27.3
LongVA-7B 14.7 2.8 5.0 0.0 10.3 0.7 0.0 6.6 2.8 0.0 49.7
Qwen2.5VL-7B 13.6 27.6 17.7 5.0 11.2 18.9 1.9 19.4 18.3 3.4 43.2
LLaVA-Next-Video-34B 12.3 0.8 17.4 0.0 11.8 0.9 0.0 6.3 9.2 0.0 48.6
LongVU-7B 12.2 3.2 1.3 0.3 4.4 2.1 0.5 3.8 1.7 0.4 34.2
Aria-28B 12.1 19.0 14.8 0.6 18.7 18.1 0.1 18.9 16.5 0.4 43.0
InternVL2-8B 7.4 14.5 8.3 1.2 9.5 9.1 0.5 12.0 8.7 0.9 40.2
Random Chance 9.8 9.8 9.8 1.0 9.8 9.8 1.0 9.8 9.8 1.0 25.0
4 Evaluation of VLMs
4.1 Evaluation Setup
We evaluate a total of 12 advanced VLMs, consisting of 4 proprietary models and 8 open-source models. Proprietary models include ChatGPT-4o (gpt-4o-2024-08-06) (OpenAI, 2024a), ChatGPT-4o-mini (gpt-4o-mini-2024-07-18) (OpenAI, 2024b), Gemini-1.5-pro (gemini-1.5-pro-002, Sep 2024) (Team et al., 2024), and Claude-3.5-sonnet (claude-3-5-sonnet-20241022) (Anthropic, 2024). Open-source models include LLaVA-Next-Video-7B (Zhang et al., 2024b), LLaVA-Next-Video-34B (Zhang et al., 2024b), LLaVA-OneVision-7B (Li et al., 2024a), LongVU-7B (Shen et al., 2024), Aria-28B (Li et al., 2024b), Qwen2.5VL-7B (Bai et al., 2025), LongVA-7B (Zhang et al., 2024a) and InternVL2-8B (Chen et al., 2024). Details about the input content and the chat template can be found in Appendix Sec. B.
Most models we tested do not natively support multiple video clips as input. As a workaround, we uniformly sample five frames from each clip and stacked them vertically as a single image. This design allows inputting multiple video clips without exceeding the image input limit of models, which is usually 32 frames. In our experiment, this trick yields the best performance. See Appendix Sec. C for details.
To prevent the model from using sentence indices as shortcuts, we input the full narration along with only a subset of the corresponding video clips. The narration and video clips are provided as separate, unaligned inputs—there is no sentence-to-clip pairing, which prevents the model from relying on structural alignment. The selected video clips span from the cause scene to the effect scene, with a random number of scenes added before and after this span. This setup further prevents shortcuts based on fixed answer location (i.e., the cause scene always appearing first and the effect scene last). Unless otherwise specified, a total of five extra scenes are padded across both ends.
4.2 Main Results
We present the performance of various VLMs on the 1-needle questions, Visual Grounding (VG) 2-needle questions, and Image Description (ID) 2-needle questions in Tab. 3. We report the accuracy on each type of questions, i.e., the proportion of correctly answered questions. For VG 2-needle questions, we separately compute the accuracy for Part 1, Part 2, and both parts answered correctly. In addition, we evaluate each model using both the original (forward) and reversed video input order, and report the average as the final result on VG 2-needle questions. The motivation of this design can be found in Sec. 4.3. We demonstrate the details of random chance calculation in Appendix Sec. B.
Figure 3: (a) The performance of models on VG 2-needle questions (Avg, Part 2) and 1-needle questions. The size of the dots indicates the model size. The stars indicate proprietary models. (b) The 2-needle visual grounding performance (Avg) on questions with different needle distances.
Two-needle Questions Are More Challenging Than One-needle.
Most models perform worse on Part 2 of VG 2-needle questions (Avg) than on 1-needle questions, despite both requiring cause clip identification. Fig. 3 (a) visualizes this pattern. Each point shows a model’s 1-needle performance (x-axis) and VG 2-needle Part 2 performance (y-axis). The diagonal line represents consistent performance across both question types, but most points fall below it, indicating that models strong on 1-needle questions often struggle with VG 2-needle questions. This highlights the deficiency of using only 1-needle questions to evaluate long-video understanding.
Open-Source Models Exhibit Weaker World Modeling Ability.
According to Tab. 3, open-source models generally perform worse than proprietary models. Specifically, proprietary models surpass all open-source models on both 1-needle and ID 2-needle questions. For VG 2-needle questions (Avg), all open-source models, except for Qwen2.5VL-7B and Aria-28B, perform at or below random levels, falling behind proprietary models. This gap may be due to insufficient human-behavior training data during the pretraining of open-source VLMs.
Performance Decreases with Increasing Needle Distance.
In Fig. 3 (b), we present the VG 2-needle question performance (Avg) for ChatGPT-4o and Gemini-1.5-pro. Performance is measured on questions with varying distances between the cause and effect events. The results show a clear decline as distance increases, especially when the evaluation aggregates answers from both parts. This phenomenon indicates that the distance between multiple needles has a significant impact when joint understanding of these needles is required. More details are provided in Appendix Sec.C.
4.3 Pathological Behaviors of VLMs
Positional Bias.
As shown in Tab. 3, the accuracy of models responses varies depending on the positions of cause and effect scenes. In VG 2-needle questions, Part 1 is grounded in the effect video clip, while Part 2 is grounded in the cause clip. In the forward-ordered clip sequence, the cause clip appears earlier, and all proprietary models show higher accuracy in Part 2 compared to Part 1; for instance, the accuracy of ChatGPT-4o exhibits an accuracy increase of 22.5%. This result appears paradoxical, since from an information-processing perspective answering the Part 2 question requires an understanding of the video clip required for the Part 1 question. Why are models more accurate when locating the cause clip required by Part 2 than Part 1?
To answer this question, we reverse the order of the input video clips (the Reverse column of Tab. 3). In this setting, Part 1 achieves much higher accuracy than Part 2. This suggests that models pay more attention to scenes appearing earlier; therefore, visual grounding for the cause clips is easier under normal playback, while reverse playback facilitates the visual grounding of the effect clips. To mitigate the influence of position bias, we evaluate model accuracy of VG 2-needle questions using the average of forward and reverse clip ordering.
Figure 4: The answer distribution of various models in the forward evaluation of visual grounding 2-needle questions. “GT” denotes ground truth. “None” means no scene number is output.
Static Output Bias.
In Tab. 3, we observe that some open-source VLMs achieve extremely low scores, sometimes even zero, on VG 2-needle questions. When examined closely, we observe that some models tend to output the same prediction regardless of the question content. We call this phenomenon the static output bias.
In Fig. 4, we compare the output distributions of well-performing and poorly-performing models on VG 2-needle questions. For example, the ground truth of Part 1 is mainly in scene 5 and beyond, as the effect scene appears later. ChatGPT-4o shows a balanced distribution, aligning with its strong performance in Tab. 3. In contrast, the three open-source models predominantly select scene 1 as the answer, demonstrating severe output bias. Instead of considering the question’s content, they produce fixed responses. To mitigate this bias, we experiment with different prompting strategies, but the models consistently produce fixed responses, indicating an inherent flaw in these models. The test prompts and output patterns are detailed in Appendix Sec. F.
Table 4: Results of models on different input settings. Numbers in parentheses indicate the absolute increase over random chance. V: Video, N: Narration, Q: Question.
Models 1-Needle Questions Image Description Questions
V+N+Q N+Q V+N+Q Q only
Proprietary Models
Gemini-1.5-pro 35.6 (+25.7) 12.3 (+2.4) 60.9 (+35.9) 39.0 (+14.0)
Claude-3-5-sonnet 26.5 (+16.7) 10.7 (+0.9) 60.5 (+35.5) 29.2 (+4.2)
ChatGPT-4o 39.2 (+29.4) 3.8 (-6.0) 59.2 (+34.2) 47.5 (+22.5)
Open-source Models
LongVA-7B 14.7 (+4.9) 11.2 (+1.4) 49.7 (+24.7) 38.4 (+13.4)
InternVL2-8B 7.4 (-2.4) 11.7 (+1.9) 40.2 (+15.2) 25.9 (+0.9)
LLaVA-Next-Video-7B 17.2 (+7.4) 0.3 (-9.5) 27.3 (+2.3) 26.9 (+1.9)
4.4 Analysis of Dataset Bias
Textual Bias.
Textual bias refers to the extent that the dataset allows models to derive the correct answer from the accompanying text alone. To verify if our visual grounding test format effectively avoids textual bias, we test various models on 1-needle questions without providing the video clips. The results are presented in the first two columns of Tab. 5. After removing video input, model performance drops significantly to around random chance (9.8%) or lower. For instance, the performance of ChatGPT-4o drops from 39.2% to 3.8%. These results indicate that our setting effectively mitigates textual bias, ensuring a reliable evaluation of multimodal understanding abilities.
Knowledge Leakage from Pretraining.
To prevent performance underestimation due to the out-of-distribution visual grounding format, our dataset introduces an image description (ID) format. However, a VLM may have memorized details of movies from pretraining, which can be used to answer these questions. To assess the impact of prior knowledge, we test different models using question-only inputs. The results are shown in the last two columns of Tab. 5. We find that ChatGPT-4o, Gemini-1.5-pro, and LongVA achieve high performance on ID questions with question-only inputs, with +22.5%, +14.0%, and +13.4% over random chance, respectively. This indicates that they can call upon prior knowledge to answer the questions. In contrast, Claude-3.5-sonnet and InternVL2-8B are less affected by prior knowledge. For almost all models tested, the performance with video and narration input is much higher than with question only, suggesting knowledge leakage alone is not sufficient for good performance on Causal2Needles.
Appendix A Dataset Statistics
Figure 5: (a) Distribution of movie themes in Causal2Needles. (b) Distribution of distance between cause and effect in Causal2Needles
The videos in Causal2Needles are collected from the SyMoN (Sun et al., 2022) and YMS (Dogan et al., 2018) datasets, comprising movie recap videos and their corresponding narrations (98 movies from SyMoN and 94 from YMS). After integration, the videos in Causal2Needles have an average duration of 438 seconds, with 58.6 clips and 1,267.2 words per movie. The movie themes in our dataset are highly diverse. To quantify this diversity, we first use ChatGPT-4o to assign multi-labels to each narrative based on 11 predefined thematic categories, and then compute the frequency of each predicted label. The distribution is shown in Fig. 5 (a). As detailed in Sec. 3, we extract 747 cause-effect pairs from the narrations and construct 2-needle questions accordingly. We further analyze the distribution of their distances, defined as the number of clips between the cause and effect events, shown in Fig. 5 (b).
Appendix B Evaluation Details
Computing Resources.
We conduct experiments using four NVIDIA GeForce RTX A6000 GPUs for all open-source models. For proprietary models, we obtain results through their publicly available APIs. The inference batch size is set to 1.
Details of VLMs.
Current video understanding VLMs primarily rely on converting videos into sequences of images for processing. For example, LLaVA-Next-Video (Zhang et al., 2024b) samples 32 frames from a video, encodes them into visual tokens using a Vision Transformer (ViT) (Dosovitskiy et al., 2020), and feeds these tokens, along with textual instructions, into Vicuna-1.5 (Zheng et al., 2023) to generate responses. Interleaved VLMs (Chen et al., 2024; Li et al., 2024b) extend input formats to support text-image sequences.
Proprietary models such as Gemini-Pro (Team et al., 2024), GPT-4o (OpenAI, 2024a), and Claude-3.5 (Anthropic, 2024) benefit from substantially greater computational resources, enabling them to process millions of image tokens—achieving dense visual understanding at up to 2 frames per second. These models also support interleaved text-image inputs, offering stronger temporal and contextual reasoning capabilities over video content.
Next, we show the input content and format we used when evaluating different models.
Proprietary LLMs.
For GPT-4o, GPT-4o-mini, Gemini-1.5-Pro-002, and Claude-Sonnet, the input consists of: (1) A series of movie scene images, including the scene containing the correct answer. Each scene is represented by five video frames stitched together vertically. (2) The task instruction, which specifies the task type—either visual grounding or image description. In the visual grounding task, the model is instructed to identify the correct scene number, whereas in the image description task, the model is asked to select the correct option. (3) The movie’s narration background. (4) The question. If the task is image description, the input will also include textual answer options. Thus, each input is structured as shown in the list below:
<image> <image> <image>
<text: Test Instruction>
<text: Movie Narration>
<text: Questions>
Open-source VLMs.
For LLaVA-Next-Video-7B, LLaVA-Next-Video-34B, LLaVA-OneVision-7B, LongVU-7B, and LongVA-7B, the input formats are the same as those used for the proprietary models. However, these models utilize different chat templates during generation. We follow the examples presented on each model’s GitHub webpage or Hugging Face repository.
LLaVA-Next-Video-7B: https://huggingface.co/llava-hf/LLaVA-NeXT-Video-7B-hf
LLaVA-Next-Video-34B: https://huggingface.co/llava-hf/LLaVA-NeXT-Video-34B-hf
LLaVA-OV-7B: https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf
LongVU-7B: https://huggingface.co/Vision-CAIR/LongVU_Qwen2_7B
LongVA-7B: https://github.com/EvolvingLMMs-Lab/LongVA
The remaining three are VLMs pretrained on interleaved image-text data. To match their input format, we insert scene numbers between video segments. However, we do not insert any text between scenes, as this could provide shortcuts by overly aligning with the narration, rather than requiring true visual understanding.
scene1<image>scene2 <image>scene3<image>
<text: Test Prompt>
<text: Movie Narration>
<text: Questions>
Random Chance Calculation.
For a 1-needle question, the model is required to select the correct scene from a sequence of given scenes (including extra padding scenes). Assuming the total number of input scenes is $N$, the random chance probability is $1/N$ (reported as a percentage). For a visual grounding 2-needle question, where the model must make two separate selections, the random chance of answering both questions correctly is $1/N^{2}$.
Since the distance between the cause and effect scenes varies, the total number of input scenes $N$ differs for each question. Therefore, we compute the average input length $\bar{N}$ over each type of questions and determine the random chance probability. For 1-needle questions, the random chance probability is 9.8%. For visual grounding 2-needle questions, the probability of correctly identifying a single part is 9.8%, while the probability of getting both correct is only 1.0%. For image description 2-needle questions with four options, the random chance is 25.0%.
Appendix C Additional Experiments
In this section, we provide a deeper analysis of several aspects: (1) the influence of the distance between cause and effect events; (2) textual grounding versus visual grounding; (3) the uniqueness of answers to Part 2 questions; and (4) non-stacked versus stacked image input settings.
The Influence of Needle Distance.
Figure 6: The 2-needle visual grounding performance of various models on questions with different needle distances.
Fig. 3 (b) in Sec. 4.2 shows that the 2-needle visual grounding performance decreases as the distance between the cause and effect events increases, based on evaluations with two proprietary models. Here, we provide additional results in Fig. 6. These results indicate that needle distance significantly impacts model performance in multi-needle questions.
Textual vs. Visual Grounding.
To better understand the mechanisms underlying the models’ performance on VG 2-needle questions, we conduct additional experiments to analyze visual and textual grounding. We introduce a new task called Sentence Grounding (SG), which requires the model to select the sentence that contains the answer from the story narration text, which are part of the input of Causal2Needles. We compare the SG performance of GPT-4o on Part 1 and Part 2 questions, and summarize the results in Fig. 7. We make the following observations.
Figure 7: Performance of ChatGPT-4o on the sentence grounding (SG) task and the visual grounding (VG) task on the 2-needle questions. In the SG tasks, the sentences are always arranged in their natural order. In the VG task, we take the average of the forward and reverse orderings.
[…]
Lastly, on SG Part 2, GPT-4o achieves 63.7% accuracy, much higher than the 30.3% on VG Part 2, suggesting the model is better at textual than visual reasoning. This large gap indicates substantial room for improvement in visual comprehension of VLMs.
Answer Uniqueness of Part 2 Questions.
We conduct an additional experiment to verify if Part 2 questions have unique answers. During causal relationship extraction, an event graph may contain multiple nodes pointing to the same node, indicating that a single event can have multiple causes. This may introduce multiple ground truth answers to Part 2 questions, as they require retrieving the cause of an event.
Our experiment is designed as follows: among the $747$ causal relationships of 2-needle questions, we identify $99$ instances where the effect sentence is not uniquely paired with a single cause sentence. For each effect sentence in them, we feed the corresponding Part 2 question and cause sentences (on average, $2.02$ candidate causes per effect) to ChatGPT-4.1. Then, it is asked to select the cause sentence that contains the answer to the Part 2 question. To avoid bias, we randomize the order of the candidate causes for each query. The results show that out of the $99$ such cases, ChatGPT-4.1 selects the wrong cause sentence in only one instance. This indicates that the answer to Part 2 question is mostly unique.
Table 5: Comparison of results between stacked image inputs and non-stacked image inputs.
Models Input Setting VG 2-Needle Questions
Forward Reverse Avg
Part 1 Part 2 Both Part 1 Part 2 Both Part 1 Part 2 Both
ChatGPT-4o non-stacked 5.0 29.4 1.4 29.7 4.1 0.9 17.4 16.8 1.2
stacked 16.7 39.2 9.4 45.4 21.2 13.4 31.1 30.2 11.4
Gemini-1.5-pro non-stacked 5.4 37.3 2.9 26.3 17.5 6.1 15.8 27.4 4.5
stacked 21.0 40.0 10.2 35.7 21.4 8.4 28.4 30.7 9.3
Non-stacked vs. Stacked Image Input.
We explain the rationale for the design choice of inputting each video clip by stacking the video frames into a single image. We represent a video clip using five uniformly sampled frames. After that, we stack these frames from top to bottom as a single image as input to the VLMs. There are two reasons for the stacked approach. First, most open-source models have a limited window size and can only process 32 frames at a time. Therefore, without stacking, processing multiple video clips would quickly exceed the window size. Second, for proprietary models that can accept more frames as input, stacking the frames can help the model better capture and comprehend the story within a video clip. As verification, we conduct additional experiments on 2-needle VG questions using ChatGPT-4o and Gemini-1.5-pro with non-stacked frames as input. For a fair comparison, we instruct the model to output the frame number containing the answer and calculate the scene number by dividing the frame number by five, as every five frames belong to a single clip. The results are detailed in Tab. 5. We can observe that stacked frames improve performance significantly, indicating that the stacked method enhances visual understanding for VLMs.
Appendix D Limitations
While our benchmark addresses several limitations of prior work, such as the lack of multi-needle evaluation on long-context video understanding and the understanding of human behaviors, there remain areas for future improvement.
First, our videos are primarily sourced from movie clips, which provide a rich set of event causal relations. Compared to real life, movies tend to over-emphasize uncommon events, such as serendipity, intricate conspiracy, or dramatic conflicts. These events could serve as effective out-of-distribution tests for models primarily trained on common events such as ego-centric videos. However, they should not be taken as necessarily representative of real-life events. Second, our current evaluation focuses on video-language models. Many of these models, particularly open-source ones, do not yet support audio inputs. Accordingly, we exclude audio from our inputs, but future extensions of the benchmark will incorporate audio to enable evaluation of more comprehensive multimodal understanding.
Appendix E Examples
We present some examples from the dataset in the Fig. 8 and Fig. 9, covering all types of questions.
Figure 8: The examples from our dataset.
Figure 9: (Continue) The examples from our dataset.
Appendix F Generation and Evaluation Prompts
F.1 Question Generation Prompts
We present prompts for generating visual grounding 2-needle questions, image description 2-needle questions, and 1-needle questions in Fig. 10, Fig. 11 and Fig. 13 respectively.
We use GPT-o1 to generate visual grounding 2-needle questions and 1-needle questions, while the large model employed for image description generation is Gemini-Pro-002. In Fig. 12, we continue using the same model to generate answer choices. At this stage, to increase the difficulty of the options, we input the cause scene as well as the two adjacent frames and prompt the model to consider generating more challenging options based on objects appearing in the remaining scenes.
Figure 10: The prompt for generating visual grounding 2-needle questions.
Figure 11: The prompt for generating image description 2-needle questions.
Figure 12: The prompt for generating options for image description 2-needle questions.
Figure 13: The prompt for generating 1-needle questions.
F.2 Question Quality Evaluation Prompts
Figure 14: The prompt used for evaluating the quality of generated questions.
We present the prompts used for evaluating the question quality (Bridge Entity, Semantic Coherence, and Readability) in Fig. 14
F.3 Cause-Effect Uniqueness Test Prompts
Figure 15: The prompt for testing the uniqueness.
We present the prompts used to test the uniqueness of the generated questions with respect to multiple candidate cause sentences in Fig. 15
F.4 Model Performance Evaluation Prompts
We present the task instruction prompts used for testing the models in Fig. 16
Figure 16: The evaluation prompts for different types of questions.
For visual grounding 2-needle questions, there are two modes: Forward playback and Reverse playback. This distinction is explicitly stated in the instruction prompt for these tasks.
F.5 Various Instruction Templates and Model Output Examples
In the Sec. 4.2 of the main paper, we observed that many open-source models exhibit severe output bias. Specifically, when testing the visual grounding 2-needle questions, these models tend to produce the same answer ("Scene 1 for part 1 and Scene 2 for part 2.") across different questions. To investigate whether this bias was caused by the prompt, we tested the models with different variations of the prompt. However, we still found that they continued to generate highly consistent responses. Below, in Fig. 17 and Fig. 18 , we list the most common outputs observed. Each type of response accounts for a significant proportion of the models’ answers across different questions. We found that although the model’s outputs varied slightly with different prompts, a significant portion of the responses remained completely identical.
Figure 17: Various test prompts, but the model fails to generate reasonable answers.
Figure 18: (Continue) Various test prompts, but the model fails to generate reasonable answers.